# Latent Adversarial Training Improves Robustness to Persistent Harmful Behaviors in LLMs

### v2

@@ A.1 RT-LAT
A.2 DPO-LAT
We now describe the DPO-LAT loss inspired by Rafailov et al. [54]. Similarly to RT-LAT, we assume that we have a paired preference dataset of harmless/harmful completions , where is the harmless result and is the harmful response. Instead of using a generic dataset of benign requests and useful completions, we instead assume is a dataset of helpful/unhelpful responses (where again is the chosen helpful response and is the rejected unhelpful one). We take from the ‘harmless’ split of Anthropic’s HH-RLHF dataset [142] and from the ‘helpful’ split.
@@ A.2 DPO-LAT
A.3 WHP-C-LAT and GA-LAT
The WHP-C-LAT and GA-LAT methods described in Section 4.3.1 and Section 4.3.2 use a toward-only adversary which optimizes for next-token cross-entropy loss on Harry Potter and the WMDP forget corpora respectively. For WHP, the model is trained as in Eldan and Russinovich [56]. For WMDP, the model uses a away loss on the forget dataset as in Mazeika et al. [21]. In both cases, we additionally include a toward loss on WikiText [145] to match Li et al. [58], and a supervised fine-tuning (SFT) loss on Alpaca [146]. While calculating the model’s toward and away losses, we keep the perturbations from the adversary. We remove these perturbations for SFT.
@@ A.3 WHP-C-LAT and GA-LAT
A.4 RMU-LAT
Here, we use the same RMU loss as used in Li et al. [58]. The adversary still optimizes for next-token cross-entropy loss on the WMDP forget corpora. In the RMU loss, when the forget loss is calculated, the adversary’s perturbation is present:
where is the length of the input tokens, and u is a randomly chosen vector from a uniform distribution between that is then normalized (and stays constant throughout training). The constants and are hyperparameter coefficients, which we set to be 6.5 and 1200 as in Li et al. [58] for Zephyr-7B.
Appendix B Jailbreaking Robustness Under Untargeted LAT
To test the advantages of targeted LAT over untargeted LAT, we compare the jailbreaking robustness of the two in Table 6. Here, during untargeted LAT, the adversary does not work to make the model comply with the jailbreak. Instead, it only works to make the model fail to output a refusal. We find that untargeted LAT results in less harm to general performance compared to targeted LAT but not refusal training. Meanwhile, untargeted lat results in comparable or slightly worse robustness in most cases compared to targeted LAT. However, for prefill and GCG attacks, untargeted LAT fares much worse than targeted LAT.
@@ Appendix B Jailbreaking Robustness Under Untargeted LAT
Appendix C Jailbreaking Robustness Under an Alternate Autograder
In Section 4.1, we evaluate jailbreak success using the StrongReject autograder [138]. However, here we also report results using the HarmBench autograder [21]. Overall, we find that the HarmBench autograder is significantly more likely to label attacks as successful, but the overall trends within results remain similar.
@@ Appendix C Jailbreaking Robustness Under an Alternate Autograder
Appendix D Backdoored Model MMLU Performance
Appendix E Low Rank Adapters and Scaled Perturbation Constraints for WHP Unlearning
In this section, we experiment with using low-rank adapters and whitened-space attacks for WHP unlearning. Typically, adversarial training methods that use projected gradient descent constrain perturbations to be within an -norm spherical ball [126]. However, for latent-space perturbations, this approach is arguably unnatural because in the latent-space, activations vary more along some directions than others. To address this, here, we test a scaling method to constrain attacks in a way that better respects the shape of the activation manifold in latent space in Section 4.3.1. We tested LAT with perturbations that are constrained to an -norm ball in whitened before they are de-whitened and added to the residual stream.
Our goal was to increase the ability of targeted LAT to operate on coherent features relating to the unlearning corpora (specifically, features that would preserve meaning but cause the model to no longer recognize the text as related). As a result, we perform principal component analysis (PCA) on the distribution of activations between Harry Potter text and the coherent genericized versions of the text produced during WHP. We optimize and constrain the perturbations in a whitened space before de-whitening them using the inverse PCA transformation matrix and then applying it to the model’s latent states. In addition, we use a low-rank adapter on all linear modules of rank 64. In our experiments, this resulted in weaker unlearning for WHP experiments but with less of a tradeoff in general capabilities. The results are shown in Table 9. However, we speculate that unlearning tasks may be especially well-suited to this type of scaling, and we leave deeper investigation to future work.
@@ Appendix E Low Rank Adapters and Scaled Perturbation Constraints for WHP Unlearning
Appendix F Tests for Robust and Competitive Unlearning in LLMs
Eldan and Russinovich [158] fine-tune Llama-2-7B-Chat [96] (Llama-2) to unlearn knowledge of the Harry Potter universe. Their method is based on fine-tuning using text that has been modified to replace domain-specific content with generic content. Throughout experiments here, we compare the WHP model from Eldan and Russinovich [56], our replications, and our replication with targeted LAT (see Section 4.3.1).
Here, we outline the methods we use to evaluate unlearning in Section 4.3.1
@@ Appendix F Tests for Robust and Competitive Unlearning in LLMs
To evaluate the model, Eldan and Russinovich [56] introduce “Familiarity” as a metric which measures the extent of Harry Potter content contained in the model’s completions of Harry Potter-related sequences as determined by an automated GPT-4 evaluation. To measure Familiarity, we follow the same method from Eldan and Russinovich [158] to evaluate a completion from the model. An evaluation prompt is formatted with the datapoint reference, prompt, and model completion, passed into GPT-4, then obtain a model Familiarity score (Figure 2), using “gpt-4-turbo-preview” at seed=42 and temperature=0, with max tokens=252. All model completions are scored in this way, and then we calculate the Familiarity metric starting a counter at 0, adding 1 for grade 3 completions, 0.2 for grade 2 completions, and 0 otherwise. Then, this total is divided by the total number of completions.
Aside from standard Familiarity evaluations as done in Eldan and Russinovich [56], we also perform four other evaluations using Familiarity, but when the model is evaluated under prompt extraction attacks.
@@
Simple jailbreaks have been successful at resurfacing knowledge that is typically not produced by LLMs (e.g., building a bomb). We test a jailbreaking prompt designed to resurface Harry Potter knowledge based on prior successful jailbreaks against Llama-2 models [144] (Figure 3).
@@
Here, we use few-shot and summary prompting. We provide the model with small amounts of general context related to Harry Potter with the goal of resurfacing existing suppressed knowledge that was not provided. We evaluate Familiarity when either a high-level summary (Figure 4) or the first 10 lines of Book 1 are included in context.
Appendix G WMDP Unlearning Details
@@
There are three layer choices that can be varied in our setup: which layer(s) of the model to put the adversary, which layers to train for RMU, and which layer to do the RMU MSE activation matching over. We kept to the same layers (trainable and RMU matching) for RMU as in Li et al. [58] – the RMU layer for the activation matching, with trainable to keep the set of hyperparameters to search over reasonably small. Applying attacks to layer requires a smaller ball radius for our random perturbations; else, we found that the adversary prevents the model trained with RMU from successfully unlearning. We also find the greatest benefit in applying attacks to the layer before the RMU activation matching layer.

### v3

@@ B.1 RT-EAT-LAT
B.2 DPO-LAT
We now describe the DPO-LAT loss inspired by Rafailov et al. (2024). Similarly to RT-EAT-LAT, we assume that we have a paired preference dataset of harmless/harmful completions , where is the harmless result and is the harmful response. Instead of using a generic dataset of benign requests and useful completions, we instead assume is a dataset of helpful/unhelpful responses (where again is the chosen helpful response and is the rejected unhelpful one). We take from the ‘harmless’ split of Anthropic’s HH-RLHF dataset (Bai et al., 2022) and from the ‘helpful’ split.
@@ B.2 DPO-LAT
B.3 WHP-C-LAT and GA-LAT
The WHP-C-LAT and GA-LAT methods described in Section˜4.3.1 and Section˜4.3.2 use a toward-only adversary which optimizes for next-token cross-entropy loss on Harry Potter and the WMDP forget corpora respectively. For WHP, the model is trained as in Eldan & Russinovich (2023). For WMDP, the model uses a away loss on the forget dataset as in Mazeika et al. (2024). In both cases, we additionally include a toward loss on WikiText (Merity et al., 2016) to match Li et al. (2024a), and a supervised fine-tuning (SFT) loss on Alpaca (Taori et al., 2023). While calculating the model’s toward and away losses, we keep the perturbations from the adversary. We remove these perturbations for SFT.
@@ B.3 WHP-C-LAT and GA-LAT
B.4 RMU-LAT
Here, we use the same RMU loss as used in Li et al. (2024a). The adversary still optimizes for next-token cross-entropy loss on the WMDP forget corpora. In the RMU loss, when the forget loss is calculated, the adversary’s perturbation is present:
where is the length of the input tokens, and u is a randomly chosen vector from a uniform distribution between that is then normalized (and stays constant throughout training). The constants and are hyperparameter coefficients, which we set to be 6.5 and 1200 as in Li et al. (2024a) for Zephyr-7B.
Appendix C Jailbreaking Robustness Under Untargeted LAT
To test the advantages of targeted LAT over untargeted LAT, we compare the jailbreaking robustness of the two in Table˜6. Here, during untargeted LAT, the adversary does not work to make the model comply with the jailbreak. Instead, it only works to make the model fail to output a refusal. We find that untargeted LAT results in less harm to general performance compared to targeted LAT but not refusal training. Meanwhile, untargeted lat results in comparable or slightly worse robustness in most cases compared to targeted LAT. However, for prefill and GCG attacks, untargeted LAT fares much worse than targeted LAT.
@@ Appendix C Jailbreaking Robustness Under Untargeted LAT
Appendix D Jailbreaking Robustness Under an Alternate Autograder
In Section˜4.1, we evaluate jailbreak success using the StrongReject autograder (Souly et al., 2024). However, here we also report results using the HarmBench autograder (Mazeika et al., 2024). Overall, we find that the HarmBench autograder is significantly more likely to label attacks as successful, but the overall trends within results remain similar.
@@ Appendix D Jailbreaking Robustness Under an Alternate Autograder
Appendix E Backdoored Model MMLU Performance
Appendix F Low Rank Adapters and Scaled Perturbation Constraints for WHP Unlearning
In this section, we experiment with using low-rank adapters and whitened-space attacks for WHP unlearning. Typically, adversarial training methods that use projected gradient descent constrain perturbations to be within an -norm spherical ball (Madry et al., 2017). However, for latent-space perturbations, this approach is arguably unnatural because in the latent-space, activations vary more along some directions than others. To address this, here, we test a scaling method to constrain attacks in a way that better respects the shape of the activation manifold in latent space in Section˜4.3.1. We tested LAT with perturbations that are constrained to an -norm ball in whitened before they are de-whitened and added to the residual stream.
Our goal was to increase the ability of targeted LAT to operate on coherent features relating to the unlearning corpora (specifically, features that would preserve meaning but cause the model to no longer recognize the text as related). As a result, we perform principal component analysis (PCA) on the distribution of activations between Harry Potter text and the coherent genericized versions of the text produced during WHP. We optimize and constrain the perturbations in a whitened space before de-whitening them using the inverse PCA transformation matrix and then applying it to the model’s latent states. In addition, we use a low-rank adapter on all linear modules of rank 64. In our experiments, this resulted in weaker unlearning for WHP experiments but with less of a tradeoff in general capabilities. The results are shown in Table˜9. However, we speculate that unlearning tasks may be especially well-suited to this type of scaling, and we leave deeper investigation to future work.
@@ Appendix F Low Rank Adapters and Scaled Perturbation Constraints for WHP Unlearning
Appendix G Tests for Robust and Competitive Unlearning in LLMs
Eldan & Russinovich (2023) fine-tune Llama-2-7B-Chat (Touvron et al., 2023) (Llama-2) to unlearn knowledge of the Harry Potter universe. Their method is based on fine-tuning using text that has been modified to replace domain-specific content with generic content. Throughout experiments here, we compare the WHP model from Eldan & Russinovich (2023), our replications, and our replication with targeted LAT (see Section˜4.3.1).
Here, we outline the methods we use to evaluate unlearning in Section˜4.3.1
@@ Appendix G Tests for Robust and Competitive Unlearning in LLMs
To evaluate the model, Eldan & Russinovich (2023) introduce “Familiarity” as a metric which measures the extent of Harry Potter content contained in the model’s completions of Harry Potter-related sequences as determined by an automated GPT-4 evaluation. To measure Familiarity, we follow the same method from Eldan & Russinovich (2023) to evaluate a completion from the model. An evaluation prompt is formatted with the datapoint reference, prompt, and model completion, passed into GPT-4, then obtain a model Familiarity score (Figure˜2), using “gpt-4-turbo-preview” at seed=42 and temperature=0, with max tokens=252. All model completions are scored in this way, and then we calculate the Familiarity metric starting a counter at 0, adding 1 for grade 3 completions, 0.2 for grade 2 completions, and 0 otherwise. Then, this total is divided by the total number of completions.
Aside from standard Familiarity evaluations as done in Eldan & Russinovich (2023), we also perform four other evaluations using Familiarity, but when the model is evaluated under prompt extraction attacks.
@@
Simple jailbreaks have been successful at resurfacing knowledge that is typically not produced by LLMs (e.g., building a bomb). We test a jailbreaking prompt designed to resurface Harry Potter knowledge based on prior successful jailbreaks against Llama-2 models (Shen et al., 2023) (Figure˜3).
@@
Here, we use few-shot and summary prompting. We provide the model with small amounts of general context related to Harry Potter with the goal of resurfacing existing suppressed knowledge that was not provided. We evaluate Familiarity when either a high-level summary (Figure˜4) or the first 10 lines of Book 1 are included in context.
Appendix H WMDP Unlearning Details
@@
There are three layer choices that can be varied in our setup: which layer(s) of the model to put the adversary, which layers to train for RMU, and which layer to do the RMU MSE activation matching over. We kept to the same layers (trainable and RMU matching) for RMU as in Li et al. (2024a) – the RMU layer for the activation matching, with trainable to keep the set of hyperparameters to search over reasonably small. Applying attacks to layer requires a smaller ball radius for our random perturbations; else, we found that the adversary prevents the model trained with RMU from successfully unlearning. We also find the greatest benefit in applying attacks to the layer before the RMU activation matching layer.